25.4 g of iodine and 12.2 g of chlorine are made to react completely to yield a mixture of ICl and ICl3. calculate the ratio of moles of ICl and ICl3.
a.  1:1
b.   1:2
c.    1:3
pls give detailed ans...

Grade:12th pass

1 Answers

Arun
5 years ago
 
It is based on principle of POAC
see given that I2 + 2CL2 gives ICL + ICL3.
Now we have molar mass of I2 =127x2=254
Molar mass of Cl2=35.5x2=71
So,,moles of I2 used in the reaction =given mass/molar mass =25.4/254=0.1 moles of I2
Moles of Cl2 used in the reaction is =14.2/71=0.2 moles of Cl2
so and therefore from the reaction we see that 1 mol of I2 react with 2 moles of Cl2 to give one mole of each ICL and ICL3 
Interpreting the same data with the same data with the given sample we can conclude that 0.2 mol of Cl2 reacts with 0.1 mol of iodine to give 0.1 mol of each ICL and ICL3
 
Regards
Arun (askIITians forum expert)